# EVTAR: End-to-End Try on with Additional Unpaired Visual Reference

### v1

@@ Abstract
We propose EVTAR, an End-to-End Virtual Try-on model with Additional Reference, that directly fits the target garment onto the person image while incorporating reference images to enhance try-on accuracy. Most existing virtual try-on approaches rely on complex inputs such as agnostic person images, human pose, densepose, or body keypoints, making them labor-intensive and impractical for real-world applications. In contrast, EVTAR adopts a two-stage training strategy, enabling simple inference with only the source image and the target garment inputs. Our model generates try-on results without masks, densepose, or segmentation maps. Moreover, EVTAR leverages additional reference images of different individuals wearing the same clothes to preserve garment texture and fine-grained details better. This mechanism is analogous to how humans consider reference models when choosing outfits, thereby simulating a more realistic and high-quality dressing effect. To support these capabilities, we enrich the training data with supplementary references and unpaired person images. We evaluate EVTAR on two widely used benchmarks and diverse tasks, and the results consistently validate the effectiveness of our approach.
Figure 1: In-the-wild try-on results generated by our end-to-end EVTAR model, trained on person and garment images from our Virtual Fitting with Reference (VFR) dataset. The first row shows results obtained by directly transferring the garment to the target person, and the second row shows results generated with additional visual references.
@@ 1 Introduction
The virtual try-on(ViTON) model aims to generate photo-realistic images of individuals wearing target clothing, with application scenarios including online retail, personalized fashion recommendation systems, and other fields. The ViTON methods can be broadly categorized into two types: Generative Adversarial Networks (GANs) [1] and diffusion models [2, 3]. Early studies typically relied on GANs [4, 5, 6]. They employed warping modules to deform clothing for alignment with the human body, followed by further fusion of the clothing and the human body to achieve visual harmony. However, such GAN-based methods generate unrealistic artifacts, especially when the clothing texture is complex or the human pose is challenging.
Recently, research based on latent diffusion models [7, 8] has gradually attracted attention in virtual try-on. Their powerful generation and editing capabilities have significantly improved clothing warping performance. Diffusion-based methods [9, 10, 11, 12] significantly address issues related to structural arrangement and texture preservation during the denoising process, which the model requires. Nevertheless, current virtual try-on technologies based on diffusion models generally still rely on additional conditions, such as clothing regions masks, garment masks, human poses, key points, and inputs of other modalities such as text prompts [8, 13, 14, 15, 16].
Figure 2: The effect of using reference images for the virtual try-on task. From left to right in the three middle subfigures are: (i) results generated without using reference images during either training or inference; (ii) results generated by a model trained with reference images but inferred without them; (iii) results generated by a model trained and inferred with reference images. Incorporating reference images consistently improves try-on detailed quality and authenticity in training and inference stages. Please zoom in for more details.
@@
Despite the remarkable progress of prior virtual try-on approaches, they are still constrained by two critical limitations that hinder their authentity of the try-on results and broader applicability: First, these approaches rely on multiple external models, such as pose estimators [17, 18, 19, 20, 21], human parse models [22, 23], segmentation models [24, 25], to process different conditions, which compromises the practicality. Moreover, in practical applications, the quality of conditional inputs—such as the cloth mask—has a substantial impact on the quality of the final try-on results; Second, many aspects of clothing, such as style, texture, and detailed design, cannot be fully perceived from the garment image alone; instead, it is more important to consider the overall appearance when the garment is worn by a model. Therefore, in real-world try-on scenarios, such as online shopping, users are typically more interested in model images rather than the garment itself. They tend to see how the target garment looks when worn on a real person, rather than relying solely on the isolated garment image as a reference. For example, as illustrated in Fig. 2, at the garment in first row, It is difficult to tell whether it is a green translucent fabric or a light green opaque one, whereas the reference image clearly reveals its green translucent material. We cannot accurately identify transparent materials or intricate designs, such as lace collars, solely from cloth images. In contrast, the reference images of human models wearing the garments reveal such details. However, existing virtual try-on methods do not support such references due to the lack of corresponding reference data in public datasets [26, 4, 27, 28, 29, 30, 31].
@@
Based on the above observation, we proposes EVTAR, an end-to-end virtual try-on framework that achieves strong performance without relying on any external models or auxiliary components, while being further enhanced by additional reference images that offer more accurate and context-aware guidance for the try-on model. First, EVTAR eliminates the need for auxiliary inputs such as segmentation masks, allowing for simple inference with only a source image and the target garment as inputs. Second, we introduce the use of images of a person clothed with the target garment as the visual references, which better reflect users’ real-world behavior when choosing clothes and enable the preservation of fine garment details that existing methods cannot achieve. Furthermore, to achieve these objectives, we construct a new dataset with supplementary reference images and use end-to-end training to train our own model, which empowers EVTAR to achieve both simplified clothing replacement and higher-quality generation results.
In summary, the main contributions of this work are as follows:
  • We propose EVTAR, a virtual try-on framework that transfers garments onto humans with or without auxiliary conditional information such as segmentation masks or dense pose. This design offers greater flexibility, significantly simplifies the model architecture, and enhances practicality for real-world applications.
  • We introduce a novel method that incorporates reference model images for virtual try-on, which greatly enhances the authenticity and the quality of the try-on results. And we construct a high-quality dataset with supplementary reference images and unpaired person images to support our method.
  • We provide a virtual try-on model that achieves state-of-the-art (SOTA) performance in both quantitative and qualitative evaluations, and also works well for in-the-wild person-cloth images, demonstrating the excellent capabilities of EVTAR.
@@ 2 Related Works
2.1 Generative Model via Flow Matching
Generative modeling has advanced rapidly, with diffusion models (DMs) [32], score-based generative models (SGMs) [33], and flow-based methods emerging as leading paradigms. Flow-based [34, 35, 36] methods have evolved to address inefficiencies in traditional continuous normalizing flows (CNFs)—which require expensive backpropagation through ODE solvers during training [37]—with a breakthrough being Flow Matching (FM) [38], which learns a time-dependent vector field to deterministically transport a simple prior distribution to the target data distribution, using a simulation-free training objective (unlike CNFs) that avoids numerical integration at training time to reduce computational overhead, and achieving comparable or superior sample quality to DMs with far fewer sampling steps (by directly parameterizing the probability flow instead of learning a stochastic reverse process [38]).
@@ 2.2 Diffusion-based Virtual Tryon
Diffusion models have advanced rapidly in recent years [2, 3], leading to a wide range of diffusion-based approaches in the virtual try-on domain [9, 10, 4]. Stable Diffusion [39, 9], with its flexible inpainting and text-guided capabilities, has become widely adopted for virtual try-on tasks. DiffusionCLIP [40] further incorporates CLIP loss to refine the generated images. DCI-VTON [41] follows a traditional two-stage pipeline, first warping clothing to align with the body and then fusing the warped garment with the person’s image. Subsequently, IDM-VTON [4] introduces a dedicated GarmentNet module that guides the garment structure and appearance, further improving visual realism. Despite these advances, diffusion-based virtual try-on methods still face notable limitations in practical application and technical completeness, leaving room for further research and optimization. In summary, existing diffusion-based virtual try-on methods either (1) rely on excessive auxiliary annotations or (2) lack support for reference images of clothed individuals. Our EVTAR model addresses these limitations by combining end-to-end diffusion training with reference-aided generation, while supporting mask-based inpaint or mask-free editing for virtual try-on.
Figure 3: The pipeline of our two-stage training strategy: (a) In the first stage, which follows a similar paradigm to mask-based try-on approaches, the model is trained on masked person images to generate person images wearing random garments for the next stage training. (b) In the second stage, the synthesized person images produced from the first stage or agnostic image, the target garment, and the additional reference images(optional) are jointly used as inputs to train an end-to-end virtual try-on model that directly fits the target cloth onto the person’s body.
@@ 3.1 Preliminary
EVTAR is built upon DiT [42], a scalable Transformer architecture for diffusion-based generation. Images are encoded into a latent space via an autoencoder [43] and then patched into tokens [44]. The diffusion process [2] operates on these tokens, with the Transformer consuming noisy tokens and predicting their denoised results.
We consider the problem of generating images conditioned on an embedding $\bm{y}$, which may encode garment images, semantic maps, human pose, or other modality-specific control signals. Let $\bm{x}$ denote the latent image representation obtained from a VAE encoder. The goal of flux.1 [45, 46] is to approximate the conditional distribution $p(\bm{x}\mid\bm{y})$ by learning a time-dependent velocity field $\bm{v}(\bm{x},\bm{y},t)$ that transports a sample from a simple prior $p_{0}(\bm{x})=\mathcal{N}(\bm{0},\bm{I})$ at $t=0$ to the data distribution $p_{\text{data}}(\bm{x}|\bm{y})$ at $t=1$. The dynamics of the conditional probability density $p(\bm{x},t|\bm{y})$ over time $t$ are governed by the continuity equation:
$\frac{\partial}{\partial t}p(\bm{x}|\bm{y},t)=-\nabla_{\bm{x}}\cdot\big(\bm{v}(\bm{x},\bm{y},t)\cdot p(\bm{x}|\bm{y},t)\big),\quad\bm{x}_{0}\sim p_{0},\ \bm{x}_{1}\sim p_{\text{data}}.$ (1)
To estimate $\bm{v}(\bm{x},\bm{y},t)$, we train a diffusion-transformer backbone to approximate the neural velocity field $\bm{v}_{\bm{\theta}}$ using the conditional flow matching objective [38, 47]:
$\mathcal{L}_{\bm{\theta}}=\mathbb{E}_{t,\bm{x}_{i},\bm{\epsilon},\bm{y}_{i}}\Big[\big\|\bm{v}_{\bm{\theta}}(\bm{x},\bm{y}_{i},t)-(\bm{x}_{i}-\bm{\epsilon})\big\|_{2}^{2}\Big],\quad\bm{x}=(1-t)\,\bm{x}_{i}+t\,\bm{\epsilon},$ (2)
@@
3.2 End To End Virtual Try-on Model with Two Stage Training
@@
Unlike previous methods that first mask the target cloth region to generate an agnostic image, our objective is to design a model that can directly fit the target cloth onto the target person, without introducing additional conditions such as densepose [17] or segmentation masks. To this end, we train a diffusion model on pairs of cloth images $\bm{c}_{i}$ and person images wearing different clothes $\bm{p}_{i}$, as illustrated in Fig. 3 (b).
@@
To train a model capable of directly taking a person’s image as input rather than relying on a segmentation mask and masked agnostic images, it is necessary to construct a training set containing unpaired person images of the form “the same person wearing different clothes,” denoted as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{\bar{p}}_{i,c_{j}}},\bm{c}_{i},\bm{p}_{i,c_{i}}]$, where the subscript $c_{j}$ of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{\bar{p}}_{i,c_{j}}$} indicates the garment other than $\bm{c}_{i}$ being worn. Unfortunately, collecting such unpaired cloth–person data pairs $\bm{\bar{p}}_{i,c_{j}},\bm{c}_{i},\bm{p}_{i,c_{i}}]$ from the real world is challenging, and these unpaired examples are not available in current open-source benchmarks sense most of them only contain data pair $\bm{c}_{i},\bm{p}_{i,c_{i}}]$. Therefore, we first synthesize unpaired person images $\bm{\bar{p}}_{i,c_{j}}$ using a mask-based try-on model, which serves as the first stage of our EVTAR training strategy. As illustrated in Fig. 3(a), we train a virtual try-on model using agnostic person images $\bm{a}_{i}$, clothing images $\bm{c}_{i}$, densepose maps $\bm{d}_{i}$, warp masks $\bm{m}i$, and other auxiliary conditions as inputs. During inference, a random unpaired garment $\bm{c}_{j}$ is selected to generate the corresponding synthesized person image $\bm{\bar{p}}_{i,c_{j}}$. To ensure the quality of the synthesized person image $\bm{\bar{p}}_{i,c_{j}}$, the agnostic–cloth pairs $\bm{a}_{i},\bm{c}_{i}]$ used for training must belong to the same garment category (e.g., if $\bm{a}_{i}$ is from the “dresses” subset, the selected garment $\bm{c}_{j}$ should also come from “dresses” rather than “upper body” or “lower body”). Otherwise, the generated person image may appear unrealistic due to mismatches between the clothing mask region and the target garment (e.g., fitting a skirt onto the upper body or a shirt onto the lower body).
3.3 Muti-input Adaptation of EVTAR

### v2

# RefVTON: person-to-person Try on with Additional Unpaired Visual Reference

@@ Abstract
We introduce RefTON, a flux-based person-to-person virtual try-on framework that enhances garment realism through unpaired visual references. Unlike conventional approaches that rely on complex auxiliary inputs such as body parsing and warped mask or require finely designed extract branches to process various input conditions, RefTON streamlines the process by directly generating try-on results from a source image and a target garment, without the need for structural guidance or auxiliary components to handle diverse inputs. Moreover, inspired by human clothing selection behavior, RefTON leverages additional reference images (the target garment worn on different individuals) to provide powerful guidance for refining texture alignment and maintaining the garment details. To enable this capability, we built a dataset containing unpaired reference images for training. Extensive experiments on public benchmarks demonstrate that RefTON achieves competitive or superior performance compared to state-of-the-art methods, while maintaining a simple and efficient person-to-person design.
Figure 1: In-the-wild try-on results generated by our RefVTON model with a p2p style, trained on person and garment images from our Virtual Fitting with Reference (VFR) dataset. The first row presents results obtained by directly transferring the garment to the target person without using any mask or human pose input. The second row shows results generated with the p2p style, incorporating additional visual references for improved accuracy and realism.
Figure 2: The effect of using reference images for the virtual try-on task. From left to right in the three middle subfigures are: (i) results generated without using reference images during either training or inference; (ii) results generated by a model trained and inferred with reference images. Incorporating reference images consistently improves the try-on quality and authenticity in both training and inference stages. Please zoom in for more details.
@@ 1 Introduction
The Virtual Try-On (ViTON) model aims to generate photo-realistic images of a person wearing target clothing, a tool crucial for applications in online retail and personalized fashion systems. ViTON methods are broadly categorized into Generative Adversarial Networks (GANs) [goodfellow2020generative] and Diffusion Models [Ho2020DenoisingDP, Rombach2021HighResolutionIS]. Early ViTON research relied on GANs [choi2021viton, han2018viton, wang2018toward], which typically employed warping modules to deform clothing for alignment with the human body, followed by fusion to achieve visual harmony. However, GAN-based approaches frequently generate unrealistic artifacts, particularly when dealing with complex clothing textures or challenging human poses. Recently, methods based on latent diffusion models (LDMs) [chen2024wear, xu2025ootdiffusion] have gained traction, significantly enhancing clothing warping and addressing structural arrangement and texture preservation during denoising [kim2024stableviton, zhu2023tryondiffusion, xu2024magicanimate, choi2024improving]. Despite these advances, current diffusion-based ViTON technologies still generally rely on extensive auxiliary conditions, such as clothing region masks, garment masks, human poses, key points, or multi-modal inputs like text prompts [xu2025ootdiffusion, kim2024promptdresser, chong2024catvton, chong2025catv2ton, yang2025omnivton].
@@
Despite the remarkable progress of prior virtual try-on approaches, they are still constrained by two critical limitations that hinder the authenticity of the try-on results and broader applicability: First, these approaches rely on multiple external models and internal modules, such as pose estimators [guler2018densepose, toshev2014deeppose, openpose, cao2017realtime, wei2016cpm], human parse models [li2020self, dong2014towards], segmentation models [kirillov2023segment, ravi2024sam2], to process different conditions, which compromises the practicality. To process diverse inputs, additional modules are integrated into the model, which consequently increases the overall framework complexity. Moreover, in practical applications, the quality of conditional inputs—such as the cloth mask—has a substantial impact on the quality of the final try-on results; Second, many aspects of clothing, such as style, texture, and detailed design, cannot be fully perceived from the garment image alone; instead, it is more important to consider the overall appearance when a model wears the garment. Therefore, in real-world try-on scenarios, such as online shopping, users are typically more interested in model images rather than the garment itself. They tend to see how the target garment looks when worn on a real person, rather than relying solely on the isolated garment image as a reference. For example, as illustrated in Fig. 2, at the garment in the first row, it is difficult to tell whether it is a green translucent fabric or a light green opaque one, whereas the reference image clearly reveals its green translucent material. We cannot accurately identify transparent materials or intricate designs, such as lace collars, solely from cloth images. In contrast, the reference images of human models wearing the garments reveal such details. However, existing virtual try-on methods do not support such references due to the lack of corresponding reference data in public datasets [morelli2022dress, choi2021viton, fang2024vivid, zheng2019virtually, shen2025imagdressing, he2024wildvidfit, meng2025hf].
@@
Based on the above observation, we propose RefTON, a flux-based person-to-person virtual try-on framework that achieves strong performance without relying on any external models or auxiliary components, while being further enhanced by additional reference images that offer more accurate and context-aware guidance for the try-on model. First, to ensure the best performance, we adopt the powerful image editing model flux-kontext as our base and apply adaptation on its position index for RoPE [su2024roformer] to make it suitable for muti-conditonal virtual try-on tasks. Similar to [chong2024catvton], RefTON eliminates the need for auxiliary inputs such as segmentation masks using a two-stage training strategy, allowing for simple inference with only a source image and the target garment as inputs. Second, we introduce the use of images of a different person clothed in the target garment as the visual references, like the model images in online shopping, which better reflect users’ real-world behavior when choosing clothes and enable the preservation of fine garment details that existing methods cannot achieve. To achieve these objectives, we propose a reference data generation pipeline, by which construct a dataset with supplementary reference images and use unpaired person-cloth sample to train our own model to utilize reference image as additional visual guidance. These improvements empower RefTON to achieve both a simplified model structure and a streamlined inference process, while simultaneously delivering superior generation results. To the best of our knowledge, RefTON is the first to support both mask-based and mask-free try-on within a single model while incorporating reference images into the pipeline.
In summary, the main contributions are as follows:
  • We propose incorporating additional reference images into the virtual try-on pipeline. This significantly enhances the authenticity and visual quality of the try-on results, achieving State-of-the-Art performance in preserving fine garment design details.
  • We designed an reference data generation framework to create the necessary reference images for both the clothing and target ground truth samples. Based on this pipeline, we built the VFR dataset upon existing benchmarks (e.g., VITON-HD, DressCode, ViViD), providing a robust new resource to improve the practicality and evaluation of virtual try-on models.
  • We present an adaptation of the Flux-Kontext I2I model with a modified positional indexing mechanism for handling multi-conditional/resolution inputs in virtual try-on. Leveraging this architecture and a two-stage training strategy, our model effectively supports both mask-based and person-to-person try-on tasks, achieving state-of-the-art results and strong generalization to in-the-wild person–clothing images.
@@ 2 Related Works
Generative modeling has advanced rapidly, with diffusion models (DMs) [sohl2015deep], score-based generative models (SGMs) [song2020score], and flow-based methods emerging as leading paradigms. Flow-based methods [dinh2016density, dinh2014nice, kingma2018glow] have evolved to address inefficiencies in traditional continuous normalizing flows (CNFs)—which require expensive backpropagation through ODE solvers during training [Chen2018NeuralOD]. A breakthrough in this area is Flow Matching (FM) [Lipman2022FlowMF], which learns a time-dependent vector field to deterministically transport a simple prior distribution to the target data distribution. Unlike CNFs, FM uses a simulation-free training objective that avoids numerical integration at training time, significantly reducing computational overhead. This approach not only achieves comparable or superior sample quality to DMs but does so with far fewer sampling steps, as it directly parameterizes the probability flow instead of learning a stochastic reverse process [Lipman2022FlowMF]. The Flux-Kontext architecture offers a powerful and flexible method for image editing. Input images are encoded into latent embeddings, which are then flattened into sequences and concatenated with Gaussian noise $\bm{\epsilon}$. Our model builds upon this architecture, leveraging its flexibility and ability to manipulate and edit images in a controlled and simplified way.
@@ 2.1 Diffusion-based Virtual Tryon
Diffusion models have advanced rapidly in recent years [Ho2020DenoisingDP, Rombach2021HighResolutionIS], leading to a wide range of diffusion-based approaches in the virtual try-on domain [kim2024stableviton, zhu2023tryondiffusion, choi2021viton]. Stable Diffusion [morelli2023ladi, kim2024stableviton], with its flexible inpainting and text-guided capabilities, has become widely adopted for virtual try-on tasks. DiffusionCLIP [kim2022diffusionclip] further incorporates CLIP loss to refine the generated images. DCI-VTON [gou2023taming] follows a traditional two-stage pipeline, first warping clothing to align with the body and then fusing the warped garment with the person’s image. Subsequently, IDM-VTON [choi2021viton] introduces a dedicated GarmentNet module that guides the garment structure and appearance, further improving visual realism. CatVTON [chong2024catvton], OmniTry [feng2025omnitry], Any2AnyTryon [guo2025any2anytryon], and OmniVTON [yang2025omnivton] investigate p2p virtual try-on methods that enable mask-free try-on. However, they either require additional conditions, such as human pose, or cannot handle mask-based and mask-free tasks simultaneously. Despite these advances, diffusion-based virtual try-on methods still face notable limitations in practical application and technical completeness, leaving room for further research and optimization.
In addition, p2p tryon methods such as TryOffDiff [velioglu2025enhancing] and ViTON-GUN [ViTON-GUN] rely on a “try-off–then–try-on” pipeline, which could easily lose garment details and accumulate errors. In contrast, our RefVTON skips the try-off stage, directly fitting the target garment onto the person. Guided by the reference image, this approach ensures superior preservation of both garment structure and material fidelity. In summary, existing diffusion-based virtual try-on methods either (1) rely on excessive auxiliary annotations or (2) lack support for reference images of clothed individuals. Our RefTON adapts the virtual try-on task to the Flux-Kontext model to address these limitations, combining person-to-person, reference-aided diffusion training while supporting both mask-based inpainting and mask-free editing for virtual try-on.
Figure 3: The pipeline of our two-stage training strategy: (a) In the first stage, which follows a similar paradigm to mask-based try-on approaches, the model is trained on masked person images to generate person images wearing random garments for the next stage training. (b) In the second stage, the synthesized person images produced in the first stage, along with the target garment and additional reference images (optional), are jointly used as inputs to train a person-to-person virtual try-on model that directly fits the target cloth onto the person’s body.
@@ 3.1 Preliminary
RefTON is built upon DiT [peebles2023scalable], a scalable Transformer architecture for diffusion-based generation. Images are encoded into a latent space via an autoencoder [kingma2013auto] and then patched into tokens [dosovitskiy2020image]. The diffusion process [Ho2020DenoisingDP] operates on these tokens, with the Transformer consuming noisy tokens and predicting their denoised results.
We consider the problem of generating images conditioned on an embedding $\bm{y}$, which may encode garment images, semantic maps, human pose, or other modality-specific control signals. Let $\bm{x}$ denote the latent image representation obtained from a VAE encoder. The goal of flux.1 [flux2024, labs2025flux1kontextflowmatching] is to approximate the conditional distribution $p(\bm{x}\mid\bm{y})$ by learning a time-dependent velocity field $\bm{v}(\bm{x},\bm{y},t)$ that transports a sample from a simple prior $p_{0}(\bm{x})=\mathcal{N}(\bm{0},\bm{I})$ at $t=0$ to the data distribution $p_{\text{data}}(\bm{x}|\bm{y})$ at $t=1$. The dynamics of the conditional probability density $p(\bm{x},t|\bm{y})$ over time $t$ are governed by the continuity equation:
$\frac{\partial}{\partial t}p(\bm{x}|\bm{y},t)$ $=-\nabla_{\bm{x}}\cdot\big(\bm{v}(\bm{x},\bm{y},t)\cdot p(\bm{x}|\bm{y},t)\big),\quad$ (1)
$\bm{x}_{0}\sim p_{0},\ \bm{x}_{1}\sim p_{\text{data}}.$
To estimate $\bm{v}(\bm{x},\bm{y},t)$, we train a diffusion-transformer backbone to approximate the neural velocity field $\bm{v}_{\bm{\theta}}$ using the conditional flow matching objective [Lipman2022FlowMF, liu2022flow]:
$\mathcal{L}_{\bm{\theta}}$ $=\mathbb{E}_{t,\bm{x}_{i},\bm{\epsilon},\bm{y}_{i}}\Big[\big\|\bm{v}_{\bm{\theta}}(\bm{x},\bm{y}_{i},t)-(\bm{x}_{i}-\bm{\epsilon})\big\|_{2}^{2}\Big],$ (2)
$\bm{x}=(1-t)\,\bm{x}_{i}+t\,\bm{\epsilon},$
@@
3.2 Person To Person Virtual Try-on Model with Two Stage Training
@@
Unlike prior methods that first mask the target clothing region to create an agnostic image, our goal is to fit the target garment onto the person directly, without relying on additional conditions such as densepose [guler2018densepose] or segmentation masks. To achieve this, we train a diffusion model on pairs of clothing images $\bm{c}_{i}$ and person images wearing different garments $\bm{\bar{p}}_{i,\bm{c}_{j}}$, as shown in Fig. 3(b). This requires constructing a training set of unpaired images of the same person in different clothes, denoted as $\bm{\bar{p}}_{i,c_{j}},\bm{c}_{i},\bm{p}_{i,c_{i}}]$, where the subscript $c_{j}$ indicates a garment other than $\bm{c}_{i}$.
@@
Unfortunately, such unpaired cloth–person data pairs $\bm{\bar{p}}_{i,c_{j}},\bm{c}_{i},\bm{p}_{i,c_{i}}]$ are not available in current open-source benchmarks and most of them only contain data pair $\bm{c}_{i},\bm{p}_{i,c_{i}}]$. Therefore, we synthesize unpaired person images $\bm{\bar{p}}_{i,c_{j}}$ using a mask-based try-on model, which serves as the first stage of our RefTON training strategy. As illustrated in Fig. 3(a), we train a virtual try-on model using agnostic person images $\bm{a}_{i}$, clothing images $\bm{c}_{i}$, densepose maps $\bm{d}_{i}$, warp masks $\bm{m}i$, and other auxiliary conditions as inputs. During inference, a random unpaired garment $\bm{c}_{j}$ is selected to generate the corresponding synthesized person image $\bm{\bar{p}}_{i,c_{j}}$. To ensure the quality of the synthesized person image $\bm{\bar{p}}_{i,c_{j}}$, the agnostic–cloth pairs $\bm{a}_{i},\bm{c}_{i}]$ used for training must belong to the same garment category (e.g., if $\bm{a}_{i}$ is from the “dresses” subset, the selected garment $\bm{c}_{j}$ should also come from “dresses” rather than “upper body” or “lower body”). Otherwise, the generated person image may appear unrealistic due to mismatches between the clothing mask region and the target garment (e.g., fitting a skirt onto the upper body or a shirt onto the lower body).
Figure 4: Adaptation of a three-channel position index: the first channel encodes different conditional inputs, while the second and third channels provide spatial positional information for adapting the resolution of the target inputs.
3.3 Multi-input Adaptation of RefTON